Draw the major organic product(s) of the following reaction. CHE NaNH / NH (0) HyC-C-CEC-H CH,CH2- CHE
Answer #1

When Terminal alkynes are treated with Sodamide/Ammonia (NaNH​​​2​ /NH​​​​​3 ) forms sodium slat of alkyne which on treatment with alkyl halides forms higher alkynes.

СН2- a MaNH2 CH - C— С=С Н OH3 сHz р: Нет — c-c=c-cң, ен. - С - С= CNo – С, он, ? -dimethyl 8- Heayne

IUPAC name of final product is 2,2 dimethyl 3-hexyne